# A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:

### v6

@@ 3 Quantitative implications
(b) Gains from a 15% GMT under $\phi=1.0$ Non-haven Haven World
@@
Notes: Panel (a) shows the effects of raising the GMT rate from $t_{M}=0$ to $0.15$ with partial coverage ($\phi=0.9$) and panel (b), from $t_{M}=0$ to $0.15$ with full coverage $\phi=1.0$. The tax rates in the initial equilibrium (Regime 0) are $0.186$ in the non-haven and $0.093$ in the haven. Results are based on the calibrated parameter values described in Table 2. The welfare levels are normalised by the MVPF, $\lambda$. The ‘haven’ here refers to the sum of the $H=40$ haven countries.
The non-haven’s revenue rises due to weaker tax competition (higher tax rates) and reduced profit shifting, while the havens’ revenues fall as the loss of shifted profits outweighs the gains from milder tax competition. The calibrated revenue gain in the non-haven (139.5 bUSD) is close to the combined gains for Europe and the US in Baraké et al., (2022, Table 1) ($67+58=125$ bEUR), and the global revenue gain (135.7 bUSD) is comparable to Hugger et al., (2024, Section 8.2) (155–192 bUSD). Since falling shifted profits reduce MNEs’ private benefits, the non-haven’s welfare rises less than their revenues. [^22] In the havens without MNE headquarters, revenues coincide with welfare.
Do the quantitative effects of introducing GMT change if the coverage is extended? The effects of $t_{M}=0.15$ with full coverage ($\phi=1.0$) yields the non-splitting equilibrium in Regime 1 (Table 3(b)). By eliminating competition for small MNEs, the extended coverage increases revenue and welfare gains worldwide by 13.6 bUSD and 8.6 bUSD, respectively. Based on the estimates by Gaul et al., (2022), these costs would amount to about 10.5 bUSD for our sample, implying a small net revenue gain (+3.1+3.1 bUSD) but a net welfare loss (1.9-1.9 bUSD) globally. [^23] Using a survey of German MNEs (see Becker et al., , 2012 for details on their characteristics), Gaul et al., (2022) estimate the average one-time expenses of implementing GMT per firm at 1.6 million EUR. Applying this to the number of small MNEs not covered by GMT in our database ($\approx$6,000; see Table A.1 in Supplemental Appendix A), the total compliance costs of newly covered MNEs are 1.6 mEUR $\times$ 6,000 $\times$ 1.1 USD/EUR $\approx$ 10.5 bUSD. Clifford et al., (2025) also estimate size-dependent compliance costs for German MNEs and similarly find that an increased coverage of the GMT would provide only limited gains for Germany.
@@ 4 Conclusion
This paper has analysed the effects of a global minimum tax (GMT) confined to large multinational enterprises (MNEs), leaving about 10% of the global tax base outside its scope. This partial-coverage GMT generates a sequence of tax competition equilibria between a non-haven and tax havens. In particular, introducing a low GMT rate that binds the havens induces them to commit to the single GMT rate. However, a further increase in the GMT rate leads first the havens, and then the non-haven, to split their tax rates for large and small MNEs.
The calibrated version of our model suggests that, upon the introduction of the current GMT rate of 15%, the most likely regime is one where the non-haven sets a uniform tax rate above the GMT for all MNEs, whereas the havens split their tax rates and undercut the GMT for small MNEs. This finding corresponds to the observation that tax havens such as Ireland and Switzerland maintain their base corporate tax rates below 15%, but top up taxes on large MNEs to the level of the GMT. In terms of welfare and tax revenues, our calibrated model predicts that the non-haven gains while the havens lose, resulting in a positive net gain globally.
@@
Appendix 1: Proof of Proposition 1
$\begin{split}&G_{n}^{1}\equiv G_{n}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}),\hskip 18.49988ptG_{h}^{1}\equiv G_{h}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}).\end{split}$
$\begin{split}&G_{n}^{2}\equiv G_{n}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}),\hskip 18.49988ptG_{h}^{2}\equiv G_{h}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}).\end{split}$
$G_{h}^{1}-G_{h}^{2}=$ (A.3)
For $t_{M}>t_{M}^{2}$, the non-haven’s single-non GMT rate, given by (A.2), is bound by the GMT and it cannot choose the unconstrained maximizing rate in response to the haven’s splitting rates. The non-haven thus chooses either a single GMT rate or a splitting tax schedule. In response, the haven must always split its tax rate; otherwise it cannot obtain positive tax revenues.
$\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\hskip 18.49988ptt_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.4)
$\begin{split}G_{n}^{3}\equiv G_{n}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}),\hskip 18.49988ptG_{h}^{3}\equiv G_{h}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}).\end{split}$
$\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\hskip 18.49988ptt_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.5)
$\begin{split}G_{n}^{4}\equiv G_{n}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}),\hskip 18.49988ptG_{h}^{4}\equiv G_{h}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}).\end{split}$
Supplemental Appendix of
‘A Global Minimum Tax for Large Firms Only:
Implications for Tax Competition’
Andreas Haufler [^1] Seminar for Economic Policy, LMU Munich, Akademiestr. 1, 80799 Munich, Germany. Email: Andreas.Haufler@econ.lmu.de.     Hayato Kato [^2] Graduate School of Economics, the University of Osaka, 1-7 Machikaneyama, Toyonaka, Osaka 560-0043, Japan. Email: hayato.kato@econ.osaka-u.ac.jp
October, 2025
Contents
Appendix A MNEs Subject to the GMT
@@
Table A.1. MNEs in the Orbis Database.
@@
Appendix B Decentralized Decision-Making by Tax Havens
In the text, we assumed that a group of tax-haven countries collectively decide whether to adopt a single GMT rate or to split their tax rate. Here, we consider an alternative, decentralized decision-making process in the first stage of our game, where each tax haven individually decides whether to split, given the decisions of the others. In the following, we focus on $t_{M}\in[t_{M}^{0},t_{M}^{2}]$ and impose two assumptions. First, countries can split their tax rates only if they are subject to the GMT. This assumption reflects the fact that preferential tax regimes for specific tax bases are not permitted under the OECD’s action plan to address base erosion and profit shifting (OECD, 2019). Second, each haven forms the same belief about how the other havens and the non-haven choose their tax rates. This allows us to focus on the problem of a single haven.
$\begin{split}&t_{n}(H_{M}-1)=\frac{2(\lambda-1)\left[t_{M}\left\{H\phi+(H_{M}-1)(1-\phi)\right\}+\delta\right]}{(\lambda-1)[\phi(H-H_{M}+1)+H+H_{M}-1]+2H\lambda},\\ &t_{h}^{2}(H_{M}-1)=\begin{cases}t_{n}(H_{M}-1)/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\end{split}$ (B.1)
The difference between the haven $h$’s payoffs in the two cases is
$\Delta G_{h}^{12}(H_{M})\equiv G_{h}^{1}(H_{M})-G_{h}^{2}(H_{M}-1)$
where
$\Theta(H_{M})\equiv AH_{M}+B,$
where
$E\equiv\delta H(2\lambda-1)(\lambda-1)>0,$
We can check $t_{M}^{0}<t_{M}^{a}<t_{M}^{b}<t_{M}^{1}$:
$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{DF}>0,$
which is indeed positive.
The discussion is summarized as follows. If $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, the non-haven chooses a single non-GMT rate, and each haven chooses a single GMT rate, as in Regime 1 in the text (Proposition 1(i)). If $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, the non-haven sets a single non-GMT rate, and each haven splits its tax rate and chooses a GMT rate for large MNEs, but a lower rate than the GMT for small MNEs, as in Regime 2 in the text (Proposition 1(ii)). If $t_{M}\in[t_{M}^{a},t_{M}^{b})$, both Regimes 1 and 2 can be equilibria and either may emerge. If $t_{M}\in[t_{M}^{a},t_{M}^{b})$ and in addition $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, the non-haven sets a single non-GMT rate ($t_{n}(H_{M}^{*})$ in (B.1)), and there are $H_{M}^{*}$ havens setting a single GMT rate $t_{M}$ and $H-H_{M}^{*}$ havens splitting their tax rates ($t_{h}^{2}(H_{M}^{*})$ in (B.1)).
A clear difference from Proposition 1 is that the range where only Regime 1 occurs is smaller ($t_{M}^{0},t_{M}^{a})\subset[t_{M}^{0},t_{M}^{1})$) when havens make their first stage decision in a decentralized manner, implying that they are less likely to commit to a single GMT rate. This is intuitive because the commitment of a haven induces the non-haven to raise its tax rate less than the simultaneous commitment of many havens.
Appendix C Welfare Effects of the Introduction of GMT
The following proposition characterizes the welfare effects of introducing a binding GMT rate, compared with the unconstrained equilibrium (Regime 0).
Proposition C.1
(GMT introduction)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text, and assume the GMT coverage rate is sufficiently high, such that $\phi\in(3/4,1)$. Compared to the unconstrained equilibrium (Regime 0), the introduction of a binding GMT rate leads to:
  • (i)
    a rise in welfare in the non-haven for all Regimes 1 to 4.
  • (ii)
    a rise in welfare in the (representative) haven for the first part of Regime 1 ($t_{M}<t_{M}^{+}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$), but a fall in welfare in the rest of Regime 1 and in Regimes 2 to 4 ($t_{M}>t_{M}^{+}$).
  • (iii)
    a rise in world welfare for all Regimes 1 to 4.
These results hold for any value of the marginal value of public funds (MVPF), $\lambda\in(1,\infty)$. In the limit as $\lambda\to\infty$, countries care only about tax revenues. Hence, qualitatively identical results apply to tax revenues.
Proof: In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
@@ Appendix C Welfare Effects of the Introduction of GMT
Appendix D Welfare Effects of Increasing the GMT Rate
@@
Proposition D.1
(Increase in the GMT rate $t_{M}$)
@@
    welfare in the non-haven increases in all regimes, except at the regime-switching point $t_{M}^{1}$, where it falls discontinuously.
@@
    welfare in the (representative) haven increases in Regime 1 as long as $t_{M}<t_{M}^{++}\equiv\frac{\delta(\lambda-1)}{2H\lambda}$, then decreases thereafter in Regime 1 and in Regime 2. It increases in Regime 3, falls discontinuously at the regime-switching point $t_{M}^{3}$, and remains unchanged in Regime 4.
@@
    welfare in the world increases in all regimes, except at the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, where it falls discontinuously.
As is true for the introduction of the GMT (Proposition C.1 in Section C of this Supplemental Appendix), the non-haven generally gains from a gradual increase in the GMT rate $t_{M}$, as this reduces profit-shifting to the havens in equilibrium. An exception is the regime-switching point $t_{M}^{1}$ where the havens being to split their tax rates and set a discretely lower tax rate for small MNEs. Small MNEs then shift more profits to the havens, hurting the non-haven.
For havens, the welfare (i.e., tax revenue) effects of an increase in the GMT rate reflect two opposing forces: a negative effect from reduced profit-shifting opportunities and a positive effect from a higher tax rate in the non-haven. These opposing forces generate a non-monotonic welfare response. In Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{1}]$), the positive effect dominates initially (when $t_{M}<t_{M}^{++}$), but as the GMT rate rises further, the negative effect prevails (when $t_{M}\geq t_{M}^{++}$). In Regime 2 ($t_{M}\in(t_{M}^{1},t_{M}^{2}]$), the negative effect dominates throughout. In Regime 3 ($t_{M}\in(t_{M}^{2},t_{M}^{3}]$), the positive effect of a higher tax rate dominates again, as the non-haven commits to a uniform GMT rate in order to raise the havens’ tax rates. This regime, where commitment induces effective tax coordination, is similar to the first part of Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{++})$). It ends at the regime-switching point $t_{M}^{3}$ at which the non-haven starts splitting its tax rate. This leads to a discontinuous fall in the non-haven’s tax rate on small MNEs, which must hurt the tax havens. In Regime 4 ($t_{M}\in(t_{M}^{3},1]$), both the non-haven and the havens apply the same GMT rate to large MNEs, preventing them from shifting profits. Consequently, changes in the GMT rate no longer affect the havens’ welfare.
The marginal effects of $t_{M}$ on world welfare are dominated by the effects on the non-haven, and are therefore generally positive. At the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, however, one of the countries begins to split its tax rate and discontinuously lowers the tax rate on small MNEs. Therefore, tax competition is tightened and world welfare falls at the margin.
@@
Proof:  In the proof of Proposition C.1 in Section C, we already showed the welfare effects of a marginal increase in the GMT rate within each regime. Here we show the welfare effects at the regime switching points.
@@
This completes the proof of Proposition D.1. ∎
Appendix E Welfare Effects of Increasing the GMT Coverage Rate

### v7

@@ 5.2.1 Decentralised decision-making by tax havens
(b) Gains from a 15% GMT with $\phi=1.0$ Non-haven Haven World
@@
Notes: Panel (a) shows the effects of raising the GMT rate from $t_{M}=0$ to $0.15$ with partial coverage ($\phi=0.9$) and panel (b), from $t_{M}=0$ to $0.15$ with full coverage $\phi=1.0$. The tax rates in the initial equilibrium (Regime 0) are $0.186$ in the non-haven and $0.093$ in the haven. The regime-switching rate between Regimes 1 and 2 is $t_{M}^{b}=9.3\%$ (see also footnote 26). Results are based on the calibrated parameter values described in Table 2. The welfare levels are normalised by the MVPF, $\lambda$. The ‘haven’ here refers to the sum of the $H=40$ haven countries.
5.2.2 Real responses of MNEs’ profits to taxes
We extend the basic model to incorporate real responses of MNEs to taxes. To keep the model simple, we assume that the pre-tax profit of an MNE decreases with the non-haven’s tax rate according to $\pi=\pi^{b}/(1+t_{n})$, where the baseline pre-tax profit, $\pi^{b}\in[\underline{\pi},\infty)$, follows a cumulative distribution function. The total pre-tax profits are then
$\Pi=\int_{\underline{\pi}}^{\infty}\frac{\pi^{b}}{1+t_{n}}dF=\frac{\Pi^{b}}{1+t_{n}},$
where $\Pi^{b}\equiv\int_{\underline{\pi}}^{\infty}\pi^{b}dF$ are the total baseline pre-tax profits. This negative effect of taxes on pre-tax profits generally reduces optimal corporate tax rates in both countries, compared to the basic model with pre-tax profits independent of taxes.
To reassess the quantitative effects of the GMT reforms, we recalibrate the parameter values because the non-haven’s tax rate in Regime 0, one of the targeted calibration moments, differs from that in the basic model. Employing the same calibration strategy as before, the new calibration yields an MVPF of $\lambda=6.0$ as shown in Table E.1 in E.2, compared to $\lambda=2.1$ in Table 2. When MNEs’ profits respond to taxes, the non-haven’s unconstrained tax rate is lower than in the absence of such responses. To match this lower rate with its data counterpart, the MVPF, which captures the relative importance of tax revenues relative to MNEs’ private benefits, must be higher. A higher MVPF raises equilibrium tax rates in both countries. In particular, the regime-switching rate between Regimes 1 and 2 is 15.9%, slightly higher than 15.6% in the basic model, above which the havens start splitting their tax rates.
The quantitative magnitudes of the GMT reforms are reported in Table E.2. in E.2 and are similar to those in Table 3. For example, the introduction of a 15% GMT rate with partial coverage increases welfare in the non-haven and the world by 108.9 and 106.3 billion USD respectively (vs. 95.6 and 92.0 billion USD in Table 3(a)), and reduces the havens’ welfare by 2.5 billion USD (vs. 3.6 billion USD in Table 3(a)).
@@ 6 Conclusion
This paper has analysed the effects of a global minimum tax (GMT) that is confined to large multinational enterprises (MNEs), thus leaving at least 10% of the global multinational tax base outside its scope. Using a simple model with profit shifting by heterogeneous MNEs, we have shown that this partial coverage of the GMT gives rise to a sequence of tax competition equilibria between a non-haven and a set of symmetric tax haven countries. In particular, introducing a low GMT rate that still binds the tax havens induces them to commit to a single GMT rate for all MNEs and thus results in higher welfare and tax revenues in both sets of countries. However, a further increase in the GMT rate leads first the tax havens, and then the non-haven, to split their tax rates for large and small MNEs, creating a conflict of interest between the two groups of countries.
The calibrated version of our model suggests that, upon the introduction of the current GMT rate of 15%, both the non-haven and the havens set a uniform tax rate (Regime 1), while introducing a 16% GMT rate leads the havens to split their tax rates and undercut the GMT for small MNEs (Regime 2). We have also checked, under alternative model specifications, that a GMT rate close to 15% results in a split tax system in the haven countries. Although it is still too early to draw definitive conclusions given the ongoing implementation of the GMT, this finding is consistent with the observation that low-tax countries such as Ireland, Hungary, and Switzerland maintain their statutory corporate tax rates below 15%, but top up the taxation of large MNEs to the level of the GMT. In terms of welfare and tax revenues, our calibrated model predicts that the non-haven gains while the tax havens lose, resulting in a positive net gain globally.
Our main objective in this paper has been to understand the implications of an incomplete coverage of the GMT. We have done so in a highly stylized model, in order to develop sharp intuitions. With the insights gained, it would be fruitful to extend our basic model to incorporate further important aspects of the GMT and the responses of MNEs. One extension would incorporate a splitting response of MNEs, whereby they restructure into smaller entities to benefit from potentially lower tax rates on small MNEs. Another extension would consider the substance-based income exclusion, i.e., additional tax deductions for MNEs that increase investment and employment. We leave these extensions for further research.
@@
Appendix A Data for Fig. 1
@@
Table A.1
MNEs in the Orbis database.
@@
Appendix B Proofs of Propositions
B.1 Proof of Proposition 1
$\begin{split}&G_{n}^{1}\equiv G_{n}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}),\qquad G_{h}^{1}\equiv G_{h}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}).\end{split}$
$\begin{split}&G_{n}^{2}\equiv G_{n}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}),\qquad G_{h}^{2}\equiv G_{h}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}).\end{split}$
$G_{h}^{1}-G_{h}^{2}=$ (A.3)
For $t_{M}>t_{M}^{2}$, the non-haven’s single-non GMT rate, given by (A.2), is bound by the GMT and it cannot choose the unconstrained maximising rate in response to the haven’s splitting rates. The non-haven thus chooses either a single GMT rate or a splitting tax schedule. In response, the haven must always split its tax rate; otherwise it cannot obtain positive tax revenues.
$\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.4)
$\begin{split}G_{n}^{3}\equiv G_{n}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}),\qquad G_{h}^{3}\equiv G_{h}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}).\end{split}$
$\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\qquad t_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.5)
$\begin{split}G_{n}^{4}\equiv G_{n}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}),\qquad G_{h}^{4}\equiv G_{h}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}).\end{split}$
B.2 Proof of Proposition 2
@@
Appendix C Equilibrium welfare and shifted profits
Fig. C.1 shows the welfare levels of the non-haven (panel (a)) and the sum of the havens (panel (b)), and the shifted profits for different GMT rates $t_{M}$.
The tax-competition equilibrium outcomes under partial GMT coverage (solid curves) are compared with those with full GMT coverage (dashed curves). The equilibrium tax rates under full GMT are as follows.
Let $G_{n}^{f}$ and $G_{h}^{f}$ be the equilibrium payoffs of the non-haven and the havens under full GMT. In Regime 0 ($t_{M}\in[0,t_{M}^{0})$), the equilibrium tax rates are identical under full and partial GMT coverage, as neither country is constrained by the GMT.
Fig. C.1. Equilibrium welfare and shifted profits.
Notes: The figure shows the equilibrium levels of welfare normalised by the marginal value of public funds (MVPF) $\lambda$ in the non-haven ($G_{n}/\lambda$) in panel (a) and the sum of the $H=40$ havens ($\sum_{h=1}^{H}G_{h}/\lambda$) in (b), and the aggregate shifted profits ($\sum_{h=1}^{H}\theta_{h}\Pi$) in (c), for different GMT rates $t_{M}$. The vertical axis is in billion USD. The solid curves in each panel are the equilibrium values under the partial GMT ($\phi\in(3/4,1)$); the dashed curves are those under the full GMT ($\phi=1$); horizontal lines are those under no GMT (Regime 0). Results are based on the calibrated parameter values described in Table 2.
Appendix D Welfare effects of gradual GMT reforms
D.1 Welfare effects of increasing the GMT rate
@@
Proposition D.1. (Increase in the GMT rate $t_{M}$)
@@
    welfare in the non-haven increases in all regimes, except at the regime-switching rate $t_{M}^{1}$, where it falls discontinuously.
@@
    welfare in the (representative) haven increases in Regime 1 as long as $t_{M}<t_{M}^{++}\equiv\frac{\delta(\lambda-1)}{2H\lambda}$, then decreases thereafter in Regime 1 and in Regime 2. It increases in Regime 3, falls discontinuously at the regime-switching rate $t_{M}^{3}$, and remains unchanged in Regime 4.
@@
    welfare in the world increases in all regimes, except at the two regime-switching rates $t_{M}^{1}$ and $t_{M}^{3}$, where it falls discontinuously.
As is true for the introduction of the GMT (Proposition 1), the non-haven generally gains from a gradual increase in the GMT rate $t_{M}$, as this reduces profit-shifting to the havens in equilibrium. An exception is the regime-switching rate $t_{M}^{1}$ where the havens being to split their tax rates and set a discretely lower tax rate for small MNEs. Small MNEs then shift more profits to the havens, hurting the non-haven.
For the havens, the welfare (i.e., tax revenue) effects of an increase in the GMT rate reflect two opposing forces: a negative effect from reduced profit-shifting opportunities and a positive effect from a higher tax rate in the non-haven. These opposing forces generate a non-monotonic welfare response. In Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{1}]$), the positive effect dominates initially (when $t_{M}<t_{M}^{++}$), but as the GMT rate rises further, the negative effect prevails (when $t_{M}\geq t_{M}^{++}$). In Regime 2 ($t_{M}\in(t_{M}^{1},t_{M}^{2}]$), the negative effect dominates throughout. In Regime 3 ($t_{M}\in(t_{M}^{2},t_{M}^{3}]$), the positive effect of a higher tax rate dominates again, as the non-haven commits to a uniform GMT rate in order to raise the havens’ tax rates. This regime, where commitment induces effective tax coordination, is similar to the first part of Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{++})$). It ends at the regime-switching rate $t_{M}^{3}$ at which the non-haven starts splitting its tax rate. This leads to a discontinuous fall in the non-haven’s tax rate on small MNEs, which must hurt the havens. In Regime 4 ($t_{M}\in(t_{M}^{3},1]$), both the non-haven and the havens apply the same GMT rate to large MNEs, preventing them from shifting profits. Consequently, changes in the GMT rate no longer affect the havens’ welfare.
The marginal effects of $t_{M}$ on world welfare are dominated by the effects on the non-haven, and are therefore generally positive. At the two regime-switching rates $t_{M}^{1}$ and $t_{M}^{3}$, however, one of the countries begins to split its tax rate and discontinuously lowers the tax rate on small MNEs. Therefore, tax competition is tightened and world welfare falls at the margin.
@@
Proof:  In B.2, we already showed the welfare effects of a marginal increase in the GMT rate within each regime. Here we show the welfare effects at the regime switching rates.
@@
This completes the proof of Proposition C.1. ∎
D.2 Welfare effects of increasing the GMT coverage rate